OpenStudy (anonymous):

Calculate the following integrals \[\Large \int_0^a dx x^{2}e^{x}\]

OpenStudy (anonymous):

In this case i need complete solution guys..

OpenStudy (anonymous):

\[ \int_0^a x^{2}e^{x}dx\]

OpenStudy (anonymous):

tomorrow i can get same question on my exam paper

OpenStudy (anonymous):

honestly i dont know how to start with it,..

OpenStudy (anonymous):

you're in calculus 1 correct?

OpenStudy (anonymous):

here in germany we name it analysis and if analysis = calculus you are right, its analysis I thema..

OpenStudy (anonymous):

but a little bit light one analysis I for computer scientist

OpenStudy (anonymous):

@UnkleRhaukus Is there anyway to solve this integral without using by parts?

OpenStudy (anonymous):

Have you heard of something called " By parts method?"

OpenStudy (anonymous):

unfortunately, not, i am complete new to integral, i have first looked at before two days ago for half hour youtube videos.. and i was not at universty when prof. lectured it =)

OpenStudy (anonymous):

Well there is certain ways of solving these types of equations that you learn later on i just don't know whether they teach it or not in analysis 1.

OpenStudy (anonymous):

ok, i am going to invest more time on integrals ;) thanks anyways guys..

OpenStudy (unklerhaukus):

Looking at my standard table of integrals, \[ [38] \qquad \int x^ne^x\text dx=x^nx^x-n\int x^{n-1}e^x \text dx\]

OpenStudy (anonymous):

Is it complete solution UnkleRhaukus ?

OpenStudy (anonymous):

this is by parts what unkle has said \[\int udv=uv-\int vdu=x^2e^x-\int e^x2xdx\\\]

OpenStudy (anonymous):

\[x^2e^x-2\int e^xxdx=x^2e^x-2[uv-\int vdu]=x^2e^x-2[xe^x-\int e^x2dx]\]

OpenStudy (anonymous):

\[x^2e^x-2xe^x+4\int e^xdx=x^2e^x-2xe^x+4e^x=e^x(x^2-2x+4)\]

OpenStudy (anonymous):

actually the last part is incorrect i took out th 2 so the du=1 and v = e^x so just get rid of one of the twos to get \ \[e^x(x^2-2x+2)\]